OpenStudy (anonymous):

A spinner is divided into 10 equal segments. Each segment is numbered 1-9 and one segment is labeled with "Free Spin." Isaiah spins the spinner ten times and gets an even number one time. What is the probability that he will get an even number on the eleventh try?

OpenStudy (anonymous):

I think you are looking at 1/11 because out of 10 spins he got an even number 1 time, so he only has a 1% chance of spinning an even number on the 11th try....but i am not completely sure

OpenStudy (anonymous):

there are four they are 2, 4, 6, and 8 ten slots all together. your probability is \[\frac{4}{10}=\frac{2}{5}=0.4\] whatever you like

Directrix (directrix):

Consider the work of @jack at http://tinyurl.com/n6xx9jn ---------- 5 Odd segments 1, 3, 5, 7, 9 5 Non-odd segments 2, 4, 6, 8, Free Spins from the past have no influence on future spins, so the chance is 5/10 = 50%